OpenStudy (anonymous):

limit of x^(sinx) as x approaches 0 from the right. how can i rewrite sin(x)ln(x) so that I get the indeterminate form of 0/0 which will allow me to apply L'hospital's rule? I tried lnx/(1/sinx) but im not getting it because ln(0) is like undefined.

OpenStudy (anonymous):

ok you got \[\sin(x)\ln(x)\] as a first step, now it is in the form \[0\times \infty\]

OpenStudy (anonymous):

usual gimmick is to rewrite with algebra as \[\frac{\ln(x)}{\frac{1}{\sin(x)}}\] or if you prefer \[\frac{\ln(x)}{\csc(x)}\] and now it is in the form \[\frac{\infty}{\infty}\]

OpenStudy (anonymous):

hmm that's what i had but I didn't know that it was in infinity/infinity form. thank you.

OpenStudy (perl):

here is a useful trick

OpenStudy (perl):

sin(x) ~ x , for x near zero

OpenStudy (perl):

therefore limit x^(sinx) -> limit x ^x , as x-> 0+

OpenStudy (perl):

and limit x^x as ->0+ is 1